### v1

# Standardization of Multi-Objective QUBOs

Abstract
Multi-objective optimization involving Quadratic Unconstrained Binary Optimization (QUBO) problems arises in various domains. A fundamental challenge in this context is the effective balancing of multiple objectives, each potentially operating on very different scales. This imbalance introduces complications such as the selection of appropriate weights when scalarizing multiple objectives into a single objective function. In this paper, we propose a novel technique for scaling QUBO objectives that uses an exact computation of the variance of each individual QUBO objective. By scaling each objective to have unit variance, we align all objectives onto a common scale, thereby allowing for more balanced solutions to be found when scalarizing the objectives with equal weights, as well as potentially assisting in the search or choice of weights during scalarization. Finally, we demonstrate its advantages through empirical evaluations on various multi-objective optimization problems. Our results are noteworthy since manually selecting scalarization weights is cumbersome, and reliable, efficient solutions are scarce.
Index Terms:
Multi-Objective Optimization, QUBO, Standardization, Quantum Annealing.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
Quantum computing is particularly well-suited for solving Quandratic Unconstrained Binary Optimization (QUBO) problems [1], which can be expressed as
$\min_{\boldsymbol{x}\in\mathcal{X}}\boldsymbol{x}^{\top}\boldsymbol{Q}\boldsymbol{x},$ (1)
where $\mathcal{X}=\{0,1\}^{n}$, $\boldsymbol{Q}$ is a symmetric real-valued $n\times n$ matrix, and $n$ denotes the number of binary variables, typically corresponding to qubits in quantum systems. Despite its seemingly simple quadratic form, the QUBO problem is NP-hard and serves as a unifying framework for a broad class of combinatorial optimization tasks. Notable application domains include chip design [2], finance [3, 4, 5], flight gate assignment [6], and various machine learning tasks [7, 8].
Although QUBO problems are widely applied in practice, finding an optimal solution is NP-hard [9], which makes exact optimization computationally infeasible for large instances. While exact algorithms exist [1], they typically exhibit exponential runtime in the worst case. As a result, practical approaches rely on (meta)heuristics that seek high-quality solutions efficiently, including simulated annealing [10], tabu search [11], and genetic algorithms [12].
(a) The difference in scale of the Pareto frontier for $\min_{\boldsymbol{x}}f(\boldsymbol{x})+g(\boldsymbol{x})$ before and after standardization. Pareto frontier was found by a brute force algorithm.
(b) The distribution of objective values of $f(\boldsymbol{x})=\boldsymbol{x}^{\top}\boldsymbol{Q}_{1}\boldsymbol{x}$ (with $n=20$) where $\boldsymbol{x}$ are sampled uniformly. Also indicated are the exact range of $f(\boldsymbol{x})$ obtained from solving it, the estimated range from its roof dual bound, as well as its exact mean and variance.
(c) The solution obtained from solving $\min_{\boldsymbol{x}}f(\boldsymbol{x})+g(\boldsymbol{x})$ exactly before and after standardization. The histograms on the x and y axes represents the distribution of $f(\boldsymbol{x})$ and $g(\boldsymbol{x})$ respectively.
Figure 1: Plots over the distribution (b) and the objective space (a,c) of the QUBO objective functions $f(\boldsymbol{x})=\boldsymbol{x}^{\top}Q_{1}\boldsymbol{x}$ and $g(\boldsymbol{x})=\boldsymbol{x}^{\top}Q_{2}\boldsymbol{x}$. The $20\times 20$ symmetric real-valued matrices $Q_{1}$ and $Q_{2}$ are randomly generated by the qubolite package [13] with random_state seeds of 00 and $1$ respectively. Additionally $Q_{2}$ is scaled up by a factor of $10$ as well.
I-A Motivation
Multi-objective optimization (MOO) involves optimizing two or more conflicting objectives simultaneously [14, 15]. This approach is essential in real-world decision-making, where trade-offs between competing goals must be carefully balanced. Unlike single-objective problems, MOO yields a set of Pareto-optimal solutions, offering decision-makers flexibility in selecting the most suitable outcome. Its importance is evident across diverse fields, including vehicle routing [16], telecommunications [17, 18], engineering design [19, 20], environmental plannig [21], and finance [22, 3].
Solution methods in MOO can be broadly categorized based on how decision-maker preferences are incorporated: no-preference, a priori, and a posteriori approaches [15, 23]. No-preference methods aim to find representative solutions without incorporating any decision-maker preferences, making them suitable when such information is unavailable or deliberately excluded. A priori methods require the decision-maker to express their preferences in advance, often through scalarization techniques that reduce the problem to a single-objective formulation [24, 25]. While effective, these methods assume that preference information is available and well-understood beforehand, which may not always be the case. In contrast, a posteriori methods first approximate or compute a diverse set of Pareto-optimal solutions, allowing the decision-maker to explore trade-offs after the optimization [26, 27, 28]. Although this reduces the need for precise prior knowledge, generating the Pareto front can be computationally intensive, especially when multiple scalarizations or high-dimensional search spaces are involved.
Given these trade-offs, we adopt a no-preference approach in this work. This choice allows us to build a simple and general foundation for demonstrating our method, without relying on strong assumptions about decision-maker preferences. Importantly, our approach can be seamlessly integrated with a priori or a posteriori techniques, which can be applied subsequently to refine the solution process based on user-defined goals or exploratory analysis.
I-B Problem Setup
Recall that in MOO, we have a set of objective functions $f_{1}(\boldsymbol{x}),f_{2}(\boldsymbol{x}),\ldots,f_{m}(\boldsymbol{x})$ we wish to optimize over simultaneously. However, in this paper, we will focus on MOO problems constructed from multiple QUBO problems—referred to as Multi-Objective Quadratic Unconstrained Binary Optimization (mQUBO) problems. Therefore, our objective functions specifically take the form
$\forall i\in\{1,\ldots,m\}:f_{i}(\boldsymbol{x})=\boldsymbol{x}^{\top}\boldsymbol{Q}^{(i)}\boldsymbol{x},$ (2)
where $\boldsymbol{x}$ is a binary vector of length $n$ and $\boldsymbol{Q}$ is a symmetric real-valued $n\times n$ matrix, similar to Equation 1.
One approach to finding a representative solution in a no-preference setting is the method of global criterion [23]. The global criterion involves minimizing the distance between the objective vector of a solution and some desirable point in the objective space. A natural choice of this desirable point is the ideal vector $\boldsymbol{z}^{\text{ideal}}$ and is constructed from the minimum value of each objective function [29],
$\min_{\boldsymbol{x}\in\mathcal{X}}\left\|f_{i}(\boldsymbol{x})-\boldsymbol{z}^{\text{ideal}}_{i}\right\|_{p},$ (3)
where $\boldsymbol{z}^{\text{ideal}}=[\min_{\boldsymbol{x}}f_{1}(\boldsymbol{x}),\ldots,\min_{\boldsymbol{x}}f_{m}(\boldsymbol{x})]$. By using the $L_{1}$-metric ($p=1$) in the global criterion, we can transform Equation 3 to a linear scalarization of our objective functions,
$\min_{\boldsymbol{x}\in\mathcal{X}}\sum_{i=1}^{m}\left|f_{i}(\boldsymbol{x})-\boldsymbol{z}^{\text{ideal}}_{i}\right|=\min_{\boldsymbol{x}\in\mathcal{X}}\left(\sum_{i=1}^{m}f_{i}(\boldsymbol{x})\right)-c,$ (4)
with $c=-\sum_{i=1}^{m}\boldsymbol{z}^{\text{ideal}}_{i}$, since the terms in the absolute values are always positive by the definition of $\boldsymbol{z}^{\text{ideal}}$. Therefore, this approach to finding a solution with no-preference is directly applicable to mQUBO as we can now just minimize the following QUBO problem:
$\min_{\boldsymbol{x}\in\mathcal{X}}\sum_{i=1}^{m}f_{i}(\boldsymbol{x})=\min_{\boldsymbol{x}\in\mathcal{X}}\boldsymbol{x}^{\top}\Biggl{(}\sum_{i=1}^{m}\boldsymbol{Q}^{(i)}\Biggr{)}\boldsymbol{x}=\min_{\boldsymbol{x}\in\mathcal{X}}\boldsymbol{x}^{\top}\tilde{\boldsymbol{Q}}\boldsymbol{x}$ (5)
to find a Pareto optimal solution.
However, in general, scalarization methods for MOO can be hampered by objective functions with differing scales as this may lead to some subset of our objective functions being over-favored during optimization. A common approach to ensure that that the objective functions are in similar scales is to normalize each objective function by dividing them with their respective ranges [30, 31, 15, 23],
$\tilde{f}_{i}(\boldsymbol{x})=\frac{f_{i}(\boldsymbol{x})-\boldsymbol{z}^{\text{ideal}}_{i}}{\boldsymbol{z}^{\text{nadir}}_{i}-\boldsymbol{z}^{\text{ideal}}_{i}},$ (6)
where $\boldsymbol{z}^{\text{nadir}}=[\max_{\boldsymbol{x}}f_{1}(\boldsymbol{x}),\ldots,\max_{\boldsymbol{x}}f_{m}(\boldsymbol{x})]$. However getting the exact range of a general QUBO objective function is equivalent to solving it exactly, and therefore impractical in the case of mQUBO problems. It is possible to estimate a lower and upper bound on the objective value of a QUBO problem by using relaxation techniques, e.g., by calculating the roof dual bound [^1] See Section III for a short description of the roof dual bound. of the problem [32] or by using semi-definite programming [33]. However, as we can see in Figure 1b, these bounds might not be tight, and therefore might end up causing the objective function to be under-weighted instead.
I-C Contribution
Therefore, we propose the use of standardization as a means to ensure that the objective functions in a Multi-Objective Quadratic Unconstrained Binary Optimization (mQUBO)are within the same scale instead; the goal of which is to encourage the solution of the scalarization in Equation 5 to be closer to the ideal point $\boldsymbol{z}^{\text{ideal}}$. See Figure 1c for an example.
Our main contributions are
  • a closed-form expression of the mean and variance for a QUBO problem’s objective values,
  • an algorithm to compute said variance with a computational complexity of $\mathcal{O}(n^{3})$ where $n$ is the number of binary variables $|\boldsymbol{x}|$,
  • and finally experiments showing the ability for standardization to result in more balanced solutions in the no-preference setting compared to using the original un-scaled objectives, or with normalization using roof dual bound estimated ranges.
The details of this approach, the algorithm, and their derivation can be found in Section II. Then in Section III, we compare the no-preference solution found by our approach with the solutions found when using the roof dual bound for normalizing QUBO objective functions, as well when no scaling approaches are used. We finally conclude this paper in Section IV with some closing thoughts and ideas for future work.
II Standardizing QUBO Objective Functions
Standardization is a commonly used method in statistics to re-scale a random variable such that it has a mean of 00 and a standard deviation of $1$. To apply it to our problem, first assume that the solution vector $\boldsymbol{x}$ can be modeled as being sampled from some probability distribution $\mathbb{P}_{\boldsymbol{X}}:\mathcal{X}\to\mathbb{R}_{[0,1]}$, where $\boldsymbol{X}$ is the random vector of binary random variables. We can then standardize the objective functions in our mQUBO problem with respect to $\mathbb{P}_{\boldsymbol{X}}$,
$\operatorname*{arg\,min}_{\boldsymbol{x}\in\mathcal{X}}\sum_{i=1}^{m}\frac{f_{i}(\boldsymbol{x})-\mathbb{E}_{\boldsymbol{X}}[f_{i}(\boldsymbol{X})]}{\sigma_{\boldsymbol{X}}(f_{i}(\boldsymbol{X}))}$ (7)
$=\operatorname*{arg\,min}_{\boldsymbol{x}\in\mathcal{X}}\Biggl{(}\sum_{i=1}^{m}\frac{f_{i}(\boldsymbol{x})}{\sigma_{\boldsymbol{X}}(f_{i}(\boldsymbol{X}))}\Biggr{)}-c.$
where $c=\sum_{i=1}^{m}\mathbb{E}_{\boldsymbol{X}}[f_{i}(\boldsymbol{X})]$ is a constant.
Therefore, the standardization approach only requires dividing the QUBO matrices by the standard deviation $\sigma_{\boldsymbol{X}}(f_{i}(\boldsymbol{X}))$ of $f_{i}(\boldsymbol{x})=\boldsymbol{x}^{\top}\boldsymbol{Q}^{(i)}\boldsymbol{x}$ with respect to some distribution $\mathbb{P}_{\boldsymbol{X}}$. Two questions arise: what is $\mathbb{P}_{\boldsymbol{X}}$ and how is $\sigma_{\boldsymbol{X}}(f_{i}(\boldsymbol{X}))$ computed?
Which distribution $\mathbb{P}_{\boldsymbol{X}}$ is correct for a specific scenario depends on various unknown factors, e.g., the actual mQUBO matrices, constraints, the underlying combinatorial problems, and so on. In the presence of this high degree of uncertainty, we choose to maximize the entropy of the underlying distribution. That is, we assume $\mathbb{P}_{\boldsymbol{X}}$ is a uniform distribution over $\mathcal{X}=\{0,1\}^{n}$.
There are two interpretations on what a uniform $\mathbb{P}_{\boldsymbol{X}}$ means. The first is that each candidate solution in $\mathcal{X}$ has a probability of $2^{-n}$ to be sampled. The second interpretation is that all random variables in $\boldsymbol{X}$ are independent Bernoulli variables distributed with a success probability of $p=0.5$. To start off this section, we will use the first interpretation to derive a closed-from expression of the mean and variance for our objective functions.
Theorem 1 (Mean and variance of QUBO with uniform $\boldsymbol{X}$)
Assume we have the objective function $f(\boldsymbol{x})=\boldsymbol{x}^{\top}\boldsymbol{Q}\boldsymbol{x}$ where $\boldsymbol{x}$ is a binary vector of length $n$ and $\boldsymbol{Q}$ is an $n\times n$ symmetric real-valued matrix. Furthermore, let $\mathbb{P}_{\boldsymbol{X}}$ be the uniform distribution over the space of all possible $n$-length binary vectors $\mathcal{X}$, $\forall\boldsymbol{x}\in\mathcal{X}:\mathbb{P}_{\boldsymbol{X}}(\boldsymbol{x})=2^{-n}$. Then the first two moments of the objective function can be computed in closed-form in time $\mathcal{O}(n^{4})$ via
$\mathbb{E}[f(\boldsymbol{X})]=\frac{1}{2}\sum_{i=1}^{n}\boldsymbol{Q}_{i,i}+\frac{1}{4}\sum_{i=1}^{n}\sum_{j\not=i}^{n}\boldsymbol{Q}_{i,j}$ (8)
and
$\mathbb{E}[f(\boldsymbol{X})^{2}]=\sum_{i,j,k,l=1}^{n}\boldsymbol{Q}_{i,j}\boldsymbol{Q}_{k,l}2^{-|\{i,j,k,l\}|}.$ (9)
Proof:
Throughout this proof, any sum over the indices $i,j,k,l$ are implied to be from $1$ to $n$ unless indicated otherwise. We shall now start by deriving the mean of $f(\boldsymbol{X})$:
$\mathbb{E}[f(\boldsymbol{X})]$ $=\sum_{\boldsymbol{x}\in\mathcal{X}}\mathbb{P}(\boldsymbol{x})\;\boldsymbol{x}^{\top}\boldsymbol{Q}\boldsymbol{x}=\sum_{\boldsymbol{x}\in\mathcal{X}}\mathbb{P}(\boldsymbol{x})\sum_{i=1}^{n}\sum_{j=1}^{n}\boldsymbol{Q}_{i,j}\boldsymbol{x}_{i}\boldsymbol{x}_{j}$
$=2^{-n}\sum_{i=1}^{n}\sum_{j=1}^{n}\boldsymbol{Q}_{i,j}\sum_{\boldsymbol{x}\in\mathcal{X}}\boldsymbol{x}_{i}\boldsymbol{x}_{j}$
$=2^{-n}\sum_{i=1}^{n}\sum_{j=1}^{n}\boldsymbol{Q}_{i,j}\times\begin{cases}2^{n-1}&\text{ if }i=j\\ 2^{n-2}&\text{otherwise}\end{cases}$
=$\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{\boldsymbol{Q}_{i,j}}{2^{|\{i,j\}|}}=\frac{1}{2}\sum_{i=1}^{n}\boldsymbol{Q}_{i,i}+\frac{1}{4}\sum_{i=1}^{n}\sum_{j\not=i}^{n}\boldsymbol{Q}_{i,j}$
where the result of the sum $\sum_{\boldsymbol{x}\in\mathcal{X}}\boldsymbol{x}_{i}\boldsymbol{x}_{j}$ depends on the number of dimensions—i.e. variables—in $\boldsymbol{X}$ that are fixed. We can then use this trick in a similar way to compute the second moment of $f(\boldsymbol{X})$ and therefore the variance,
$\mathbb{E}[f(\boldsymbol{X})^{2}]$ $=\sum_{\boldsymbol{x}}\mathbb{P}(\boldsymbol{x})\left(\sum_{i=1}^{n}\sum_{j=1}^{n}\boldsymbol{Q}_{i,j}\boldsymbol{x}_{i}\boldsymbol{x}_{j}\right)^{2}$
$=\sum_{\boldsymbol{x}}\mathbb{P}(\boldsymbol{x})\sum_{i,j,k,l=1}^{n}\boldsymbol{Q}_{i,j}\boldsymbol{Q}_{k,l}\boldsymbol{x}_{i}\boldsymbol{x}_{j}\boldsymbol{x}_{k}\boldsymbol{x}_{l}$
$=\sum_{i,j,k,l=1}^{n}\boldsymbol{Q}_{i,j}\boldsymbol{Q}_{k,l}2^{-|\{i,j,k,l\}|},$
thus proving the equality in Equation 9. ∎
While the variance derivation in Theorem 1 is succinct, direct computation of the sum in Equation 9 has a time complexity of $\mathcal{O}(n^{4})$, which becomes computationally expensive for large $n$. To address this, Theorem 2 derives a closed-form expression for the variance of $f(\boldsymbol{X})$ with a reduced computational complexity of $\mathcal{O}(n^{3})$, which can lead to significant speedups for large $n$.
Theorem 2 (Variance of QUBO with independent Bernoulli distributions $\mathcal{B}(0.5)$)
Assume we have the objective function $f(\boldsymbol{x})=\boldsymbol{x}^{\top}\boldsymbol{Q}\boldsymbol{x}$ where $\boldsymbol{x}$ is a binary vector of length $n$ and $\boldsymbol{Q}$ is a symmetric real-valued matrix. Then further assume that $\boldsymbol{X}$ is a vector of independent Bernoulli random variables with success probability $p=0.5$, $\forall i\in\{1,2,\ldots,n\}:\boldsymbol{X}_{i}\sim\mathcal{B}(0.5)$. The variance of $f(\boldsymbol{X})$ is then:
$\operatorname{Var}(f(\boldsymbol{X}))$ (10)
$=\sum_{i}\frac{1}{4}\boldsymbol{Q}_{i,i}^{2}+\frac{1}{8}\Biggl{(}\sum_{k\neq i}\boldsymbol{Q}_{i,i}\cdot[\boldsymbol{Q}_{k,i}+\boldsymbol{Q}_{i,k}]\Biggr{)}$
$\quad+\sum_{i}\sum_{j\neq i}\boldsymbol{Q}_{i,j}\Biggl{(}\frac{3}{16}\Bigl{(}\boldsymbol{Q}_{i,j}+\boldsymbol{Q}_{j,i}\Bigr{)}+\frac{1}{8}\Bigl{(}\boldsymbol{Q}_{i,i}+\boldsymbol{Q}_{j,j}\Bigr{)}\Biggr{)}$
$\quad+\sum_{i}\sum_{j\neq i}\boldsymbol{Q}_{i,j}\sum_{k\neq\{i,j\}}\frac{1}{16}\Bigl{(}\boldsymbol{Q}_{k,i}+\boldsymbol{Q}_{k,j}+\boldsymbol{Q}_{i,k}+\boldsymbol{Q}_{j,k}\Bigr{)}.$
Proof:
First notice that the product of any Bernoulli variable with itself results in the same random variable, $\boldsymbol{X}_{i}\cdot\boldsymbol{X}_{i}=\boldsymbol{X}_{i}$. Furthermore, the product of two idependent Bernoulli random variables $\boldsymbol{X}_{i},\boldsymbol{X}_{j}\sim\mathcal{B}(p)$ results in the new Bernoulli random variable $\boldsymbol{X}_{i}\cdot\boldsymbol{X}_{j}\sim\mathcal{B}(p^{2})$. Then recall the following equivalence between the variance of a sum of random variables and sum of their pairwise covariance, also known as Bienaymé’s identity,
$\operatorname{Var}(f(\boldsymbol{X}))=$ $\operatorname{Var}\biggl{(}\sum_{i,j}\boldsymbol{X}_{i}\boldsymbol{X}_{j}\boldsymbol{Q}_{i,j}\biggr{)}$
$=$ $\sum_{i,j}\sum_{k,l}\boldsymbol{Q}_{i,j}\boldsymbol{Q}_{k,l}\operatorname{Cov}(\boldsymbol{X}_{i}\boldsymbol{X}_{j},\boldsymbol{X}_{k}\boldsymbol{X}_{l}).$
Note that for $\operatorname{Cov}(\boldsymbol{X}_{i}\boldsymbol{X}_{j},\boldsymbol{X}_{k}\boldsymbol{X}_{l})$ to be non-zero, $\boldsymbol{X}_{i}\boldsymbol{X}_{j}$ and $\boldsymbol{X}_{k}\boldsymbol{X}_{l}$ needs to share a common variable since if they do not,
$\operatorname{Cov}(\boldsymbol{X}_{i}\boldsymbol{X}_{j},\boldsymbol{X}_{k}\boldsymbol{X}_{j})$
$=\mathbb{E}[\boldsymbol{X}_{i}\boldsymbol{X}_{j}\boldsymbol{X}_{k}\boldsymbol{X}_{j}]-\mathbb{E}[\boldsymbol{X}_{i}\boldsymbol{X}_{j}]\mathbb{E}[\boldsymbol{X}_{k}\boldsymbol{X}_{j}]$
$=2^{-|\{i,j\}|-|\{k,l\}|}-2^{-|\{i,j\}|}\cdot 2^{-|\{k,l\}|}$
$=0.$
Therefore, one approach we can use to compute the variance is to enumerate all the cases when $\operatorname{Cov}(\boldsymbol{x}_{i}\boldsymbol{x}_{j},\boldsymbol{x}_{k}\boldsymbol{x}_{l})$ is non-zero.
  1. 1.
    When $i=j$, and
    1. (a)
      $k=l=i$, the covariance is $1/4$.
    2. (b)
      $k=i$ but $l\neq i$, the covariance is $1/8$.
    3. (c)
      $l=i$ but $k\neq i$, the covariance is $1/8$.
  2. 2.
    When $i\neq j$, and
    1. (a)
      $k=i$ and $l=j$, the covariance is $3/16$.
    2. (b)
      $k=j$ and $l=i$, the covariance is $3/16$.
    3. (c)
      $k=i$ and $l=i$, the covariance is $1/8$.
    4. (d)
      $k=j$ and $l=j$, the covariance is $1/8$.
  3. 3.
    When $i\neq j$, and
    1. (a)
      $k=i$ and $l\notin\{i,j\}$, the covariance is $1/16$.
    2. (b)
      $k=j$ and $l\notin\{i,j\}$, the covariance is $1/16$.
    3. (c)
      $l=i$ and $k\notin\{i,j\}$, the covariance is $1/16$.
    4. (d)
      $l=j$ and $k\notin\{i,j\}$, the covariance is $1/16$.
As a result, we can now compute the variance of $f(\boldsymbol{X})$ by just taking the sum over the indices where we know the covariance term is non-zero and grouping terms with the same covariance together, resulting in Equation 10. ∎
Furthermore, when implementing the expression in Equation 10, it is also possible to further simplify the terms in the sum. This results in a concise algorithm for the computation of the variance for any QUBO objective function—as seen in Figure 2—with a computational complexity of $\mathcal{O}(n^{3})$.
Input: QUBO matrix $Q$
Output: Variance $v$
$n\leftarrow$ number of rows in $Q$;
$v\leftarrow 0$;
for $i\leftarrow 1$ to $n$ do
       for $j\leftarrow 1$ to $n$ do
             if $i\neq j$ then
                   $v\leftarrow v+\boldsymbol{Q}_{i,j}(\boldsymbol{Q}_{i,j}+\boldsymbol{Q}_{j,i})/16$;
             end if
            for $k\leftarrow 1$ to $n$ do
                   $v\leftarrow v+\boldsymbol{Q}_{i,j}(\boldsymbol{Q}_{k,i}+\boldsymbol{Q}_{k,j}+\boldsymbol{Q}_{i,k}+\boldsymbol{Q}_{j,k})/16$;
             end for
       end for
end for
return $v$;
Algorithm 1 Fast QUBO Variance in $\mathcal{O}(n^{3})$
Figure 2: Algorithm for computing the variance of a QUBO objective function.
III Experiment
We now wish to showcase the applicability of our approach of using standardization for scaling the objective functions in mQUBO problems. Recall from Section I-B we argued that the main benefit of scaling the objective functions is to discourage any single objective function from being over-prioritized in the final scalarized compound objective function. Therefore in this section, we will compare the solutions we get from different scaling approaches when solving the equal weighted linear scalarization of our objective functions. Specifically we will compare,
  1. 1.
    a scalarization where none of the objective functions are scaled,
  2. 2.
    an approach that normalizes the objectives by estimating the range of each QUBO objective function via their roof dual bound, and
  3. 3.
    our approach of scaling via standardization.
The roof dual bound [32] is a relaxation technique for QUBO problems that provides a tight lower bound by solving a related continuous optimization problem. It leverages roof duality theory to identify variables that can be fixed to binary values—known as persistent variables—based on the structure of the quadratic objective. This is often achieved through a max-flow formulation. Variables with strong persistency are guaranteed to have the same values in any optimal solution, enabling effective preprocessing and problem size reduction with a worst-case runtime of $\mathcal{O}(n^{3})$. In our experiments, we specifically use the implementation found in qubolite [13].
III-A QUBO Problems
We will now lay out the mQUBO problems and methodology we will use to compare the discussed scaling methods. To start we first define 4 different QUBO problems and their respective objective functions for use in our experiments. The following max cut and subset sum problems [34] are all defined on the same set of vertices, $V(\boldsymbol{G})$:
  1. 1.
    MC$[0,1]$: A max cut problem where the weights on the graph $\boldsymbol{G}$ are sampled from the beta distribution, $\forall(i,j)\in E(\boldsymbol{G})$:
    $\boldsymbol{w}_{i,j}^{(P)}\sim\operatorname{Beta}(0.2,0.8)$
  2. 2.
    MC$\{0,1\}$: A max cut problem defined on the complete graph over the vertices of $\boldsymbol{G}$, $K_{\boldsymbol{G}}$. The weights of each edge on the complete graph are sampled from a Bernoulli distribution with equal probability of sampling 00 or $1$. However, the weight of edges that already exist in $E(\boldsymbol{G})$ are always $1$:
    $\forall(i,j)\in E(K_{\boldsymbol{G}}):\begin{cases}\boldsymbol{w}_{i,j}^{(B)}=1&(i,j)\in E(\boldsymbol{G})\\ \boldsymbol{w}_{i,j}^{(B)}\sim\mathcal{B}(0.5)&\text{otherwise}\\ \end{cases}$
  3. 3.
    MC$[1,5]$: A max cut problem defined on the complete graph over the vertices of $\boldsymbol{G}$, $K_{\boldsymbol{G}}$. The weights of each edge on the complete graph are sampled from a discrete uniform distribution from $1$ to $5$, $\mathcal{U}\{1,5\}$. However, the weight of edges that already exist in $E(\boldsymbol{G})$ are always $5$:
    $\forall(i,j)\in E(K_{\boldsymbol{G}}):\begin{cases}\boldsymbol{w}_{i,j}^{(Z)}=5&(i,j)\in E(\boldsymbol{G})\\ \boldsymbol{w}_{i,j}^{(Z)}\sim\mathcal{U}\{1,5\}&\text{otherwise}\\ \end{cases}$
  4. 4.
    SubSum: A subset sum problem defined on the vertices of $\boldsymbol{G}$, $V(\boldsymbol{G})$, where the weights of each vertex is the number of neighbors it has,
    $\forall v\in V(\boldsymbol{G}):\boldsymbol{w}_{v}^{(S)}=|N_{\boldsymbol{G}}(v)|.$
    We also use a quarter of the total vertex weights as the target sum for the problem,
    $\tau=\frac{1}{4}\sum_{v\in V(\boldsymbol{G})}\boldsymbol{w}_{v}^{(S)}.$
The max-cut problems MC$[0,1]$, MC$\{0,1\}$, and MC$[1,5]$ have objective functions with the following form [35]:
$f(\boldsymbol{x})=\sum_{i=1}^{n}\sum_{j=i+1}^{n}2\boldsymbol{w}_{i,j}\boldsymbol{x}_{i}\boldsymbol{x}_{j}-\sum_{i=1}^{n}\boldsymbol{x}_{i}\Biggl{(}\sum_{j=1,j\neq i}^{n}\boldsymbol{w}_{i,j}\Biggr{)}$
where $\boldsymbol{w}\in\{\boldsymbol{w}^{(P)},\boldsymbol{w}^{(B)},\boldsymbol{w}^{(Z)}\}$. On the other hand, the subset sum problem has the following objective function [36]:
$f_{S}(\boldsymbol{x})=\sum_{ij}\boldsymbol{w}_{i}^{(S)}\boldsymbol{w}_{j}^{(S)}\boldsymbol{x}_{i}\boldsymbol{x}_{j}-2\tau\sum_{i}\boldsymbol{w}_{i}^{(S)}\boldsymbol{x}_{i}.$
These QUBO objectives can be motivated by the real-life problem of marketing a multiplayer video game and selecting a subset of influencers to be part of the marketing campaign. Here MC$[0,1]$ represents the friendship between influencers and how likely they are to get another influencer to play the game if they are part of the campaign. MC$\{0,1\}$ and MC$[1,5]$ on the other hand indicates if any two influencer share a common language and how compatible their timezones are respectively. And finally SubSum represents the budget we want to use up in the campaign.
III-B Experimental setup
For our experiments we instantiate instances of our QUBO problems with a random Barabasi-Albert graph with $1000$ nodes and a connectivity of $2$ using the NetworkX library [37] with a seed of $1$. The distributions used for problem generation are initialized using numpy [38] with seeds of $1$ as well.
We then choose all possible combinations of the four QUBO problems leading to $\binom{4}{2}+\binom{4}{3}+\binom{4}{4}=11$ different combinations, and therefore mQUBO problems for our experiments. For each mQUBO problem, we use the 3 different approaches for objective function scaling to create 3 different scalarizations of the mQUBO problem, resulting in a total of $33$ different scalarized QUBO problems to solve. We also define the operator $\boldsymbol{F}$ for each of the $11$ mQUBO problems such that $\boldsymbol{F}(\boldsymbol{x})$ returns an $m$-length vector in objective space, where $m$ is the number of objectives.
These scalarizations are then optimized using the digital annealer from Fraunhofer IAIS [39]. The digital annealer is given a time limit of $2$ seconds per run to solve each given scalarized problem. We then run the digital annealer $20$ times for each scalarized problem, before returning a set of solutions found over those $20$ runs. We filter this set of solutions to find the set of non-dominated solutions $\boldsymbol{S}$ found by the digital annealer. A solution $\boldsymbol{x}$ is said to dominate another solution $\boldsymbol{y}$ if every element of $\boldsymbol{x}$ is less than or equal to the corresponding element of $\boldsymbol{y}$, and at least one element in $\boldsymbol{x}$ is strictly less than the corresponding element in $\boldsymbol{y}$ [23],
$(\forall i\in\mathbb{N}_{[1,n]}\;\boldsymbol{x}_{i}\leq\boldsymbol{y}_{i})\wedge(\exists j\in\mathbb{N}_{[1,n]}\;\boldsymbol{x}_{j}<\boldsymbol{y}_{j}).$ (11)
This entire process is repeated $20$ times to get $20$ sets of non-dominated solutions, $\mathcal{S}=\{\boldsymbol{S}_{1},\ldots,\boldsymbol{S}_{20}\}$.
TABLE I: Mean Hypervolume of the set of non-dominated solutions found by the digital annealer over 20 runs. Highest means for each mQUBO problem are in bold.
QUBO Problems Original Roof Dual Standardization
MC[0,1] MC{0,1} MC[1,5] SubSum Mean Std Dev Mean Std Dev Mean Std Dev
3.005e+06 2.276e-09 3.18e+06 4.787e+04 5.071e+06 1.818e+04
1.128e+07 2.484e-08 1.232e+07 3.847e+04 1.849e+07 5.265e+04
9.894e+08 1.621e+06 1.106e+09 9.095e+06 1.666e+09 6.928e+06
3.038e+04 2.44e-11 3.029e+04 5.699e-11 3.029e+04 5.699e-11
1.644e+11 4.971e+07 1.635e+11 0.0002572 1.973e+11 2.286e+08
3.786e+11 2.973e+07 3.793e+11 0.0007165 4.768e+11 1.935e+09
1.828e+10 2.17e-05 6.281e+09 6.523e+06 2.368e+10 2.322e-05
1.18e+14 1.65e+11 3.486e+14 6.299e+12 4.567e+14 2.129e+12
2.237e+14 7.501e+09 5.879e+14 1.423e+13 8.484e+14 4.755e+12
2.353e+16 50.35 7.004e+16 24.00 8.04e+16 59.53
7.934e+18 1.146e+14 6.45e+19 2.484e+16 8.099e+19 7.914e+15
III-C Performance Measure: Hypervolume Indicator
Figure 3: Example of reference points (black) used for the Hypervolume calculations—that are later averaged over—in 2D objective space over problems MC$\{0,1\}$ and MC$[0,1]$.
The measure we use to compare the solutions in $\boldsymbol{S}\in\mathcal{S}$ between the different scaling methods is to compute the hypervolume of $\boldsymbol{S}$ found from optimizing the scalarizations corresponding to each method. Hypervolume essentially measures the quality and diversity of a set of non-dominated solutions by finding the size of the objective space bounded by a reference point and dominated by the objective vector of at least one solution $\boldsymbol{F}(\boldsymbol{x}),\boldsymbol{x}\in\boldsymbol{S}$ [40]. The reference point used must be dominated by all the solutions in $\boldsymbol{S}$ in objective space.
For our experiments, we actually use a set of $10000$ uniformly generated reference points over the hypercube defined by two points, $\boldsymbol{z}^{\text{ref}}$ and $2\cdot\boldsymbol{z}^{\text{ref}}-\boldsymbol{z}^{\text{desire}}$, and average the computed hypervolume over these reference points. Here the point $\boldsymbol{z}^{\text{ref}}$ is unique to each of the 11 mQUBO problems and is the element-wise maximum in objective space over all $\boldsymbol{S}\in(\mathcal{S}_{N}\cup\mathcal{S}_{S}\cup\mathcal{S}_{O})$, where $(\mathcal{S}_{N},\mathcal{S}_{S},\mathcal{S}_{O})$ corresponds to the set of sets of solutions from optimizing the scalarization with normalization, standardization, and with no scaling respectively. On the other hand, $\boldsymbol{z}^{\text{desire}}$ represents the element-wise minimum over the same set of solutions instead. The hypercube defined between $\boldsymbol{z}^{\text{ref}}$ and $2\cdot\boldsymbol{z}^{\text{ref}}-\boldsymbol{z}^{\text{desire}}$ is then just a translated copy of the hypercube defined by the points $\boldsymbol{z}^{\text{desire}}$ and $\boldsymbol{z}^{\text{ref}}$. An example of these randomly generated reference points for a two objective problem can be found in Figure 3.
III-D Results
TABLE II: The roof dual range and standard deviation for each QUBO problem used in our experiments.
QUBO Roof Dual Range Std. Deviation
MC[0,1] 7.578e2 2.235e1
MC{0,1} 2.520e5 4.605e3
MC[1,5] 1.507e6 2.753e4
SubSum 2.680e7 4.216e5
From Table I, we can see that standardization almost always score higher on the hypervolume indicator, implying that the solutions we get from standardizing the objective functions of a mQUBO before equal weighted scalarization seems to be providing more balanced trade-offs between the different objectives.
However, for the mQUBO problem with the objectives MC$\{0,1\}$ and MC$[1,5]$, the solution found for the scalarization without any scaling performs best. That said, the hypervolume indicators for this mQUBO problem are very close to each other, with a maximum difference of $0.2962$%. We hypothesize that the reason for the similar performances is partly due to MC$\{0,1\}$ and MC$[1,5]$ being defined very similarly; they are both max-cut problems on the complete graph where edges that were already in the original graph are given the largest weight. Furthermore, from Table II, we can see that the roof dual range and standard deviation of MC$\{0,1\}$ and MC$[1,5]$ are different by just a single order of magnitude. So we believe it is a combination of both factors that led to the different approaches performing similarly.
IV Conclusion
When multiple objective functions have to be addressed simultaneously, setting approriate preferences is a notoriously hard and cumbersome process. No-preference methods aim to find representative solutions but implicitly require that all invovled objectives have a similar scale.
To this end, we proposed the use of standardization as a means to ensure a set of QUBO objective functions are within a similar scale—as opposed to normalization. The main benefit of using standardization is that it can be done exactly with a computational complexity of $\mathcal{O}(n^{3})$. We proved the computational complexity of standardization by both deriving a closed-form expression for the variance of a QUBO objective function when the binary solution vector $\boldsymbol{x}$ is sampled from a uniform distribution. We also provided a concise algorithm for computing this variance. Finally, we showed that solutions obtained from equal weighted scalarization with standardization scores higher on the hypervolume indicator on average compared to other scaling approaches. Therefore standardization provides a more balanced trade-off between the different objectives compared to just using the original objective function or normalization using a roof dual bound estimate on the range of the objectives.
One possible avenue of future work would be to theoretically analyze the effects on the optimization landscape from scaling QUBO objectives by its standard deviation. Furthermore, it would be interesting to see how this approach of standardization can affect the integration of constraints into an mQUBO problem as well. Additionally, since we’ve only considered the variance when $\boldsymbol{x}$ is sampled from a uniform distribution, an interesting direction for future research would be to derive versions of Theorem 1 for other distributions of $\boldsymbol{X}$ and studying its impact on the pareto front, especially when subspaces of $\mathcal{X}$ become infeasible due to hard constraints.
Acknowledgment
This research has been funded by the Federal Ministry of Education and Research of Germany and the state of North-Rhine Westphalia as part of the Lamarr-Institute for Machine Learning and Artificial Intelligence.